In an experiment for measuring the young's modulus (y), following readings were taken. Load =3 kg, length = 2.820 m , diameter = 0.041 cm and extension = 0.87. Determine percentage error in measurement of Y.

Grade:11

1 Answers

Ramesh V
14 years ago

Dear vishrant,

I'm not sure whether this is the correct method to analyse the error,

The relation is Y = (FL) / (A.dl) where F=mg and A = Pi.D2 /4 (take g = 10m.sec2 )

to calculate Y , given F=30 N ,dL/L=0.87 and A=1.32x10-7 m2

so Y = 261.23 MPa

consider dl , this term is a change in lenth, so error is nullified as we are taking the change in length : d(dl)=0

also change in gravity dg = 0

as Y = (4mgL) / (Pi.D2.dl)

 ln Y = ln 4 + ln m + ln g + ln L - 2.ln D - ln (dl)

dY/Y = dm/m  +dL/L + dL/L + 2.dD/D   ( as d(dl)=0 and dg =0 )

Taking the least counts for measuring load ( using pan balance range 5kgs,LC =1g)  = 1 gram =10-3kg

Taking the least counts for measuring length ( using meter scale range 1m,LC =0.1cm)  = 1 mm =10-3m

Taking the least counts for measuring diameter ( using vernier calipers ,LC =0.01cm)  = 0.01 mm =10-5m


dY/Y = dm/m  +dL/L + dL/L + 2.dD/D

         = 10-3 / 3 +10-3 / 2.82 + 2*10-5/41*10-5
         =0.0495

so (Y±dY )= (261.23 ± 12.93) MPa


or percentage error is : dY/Y x 100 = 4.95 %


-----

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch